AN INTERPOLATION THEOREM FOR HOLOMORPHIC AUTOMORPHISMS OF C^{n}

GREGERY T. BUZZARD FRANC FORSTNERIC

Abstract. We construct automorphisms of C^{n} which map certain discrete sequences one onto another with prescribed finite jet at each point, thus solving a general Mittag-Leffler interpolation problem for automorphisms. Under certain circumstances, this can be done while also approximating a given automorphism on a compact set.